An electron in a hydrogen atom in its ground state absorbs energy equal to the ionisation energy of Li+2 the wavelength of emitted electron is

Grade:11

1 Answers

Khimraj
5 years ago
Ionization energy for Li2+= 13.6 Z²/n² eV
Z=3. Thus, I.E= 122.4 eV.
Energy needed to eject electron from H in ground state=13.6 eV.
Kinetic energy of electron= Energy supplied-Energy required.
=> 122.4eV -13.6 eV.= 108.8 eV.
Now use De Broglie equation,
wavelength = hc/E
= 6.63*10-34*3*108/108.8*1.6*10-19 =  114.2 Ao
Hope it clears.